Wenbo is a monopolist in the market for peppa pig figurines within the economics department. He has the following information, where Q is the quantity of peppa pig figurines and P is the price: Market Demand: P = 300 – Q

Marginal Cost: MC = 2Q

Total Cost: TC = 2 + ?2

a. Suppose Wenbo cannot price discriminate. What is the profit maximizing price and quantity for Wenbo? What is Wenbo’s profit? What is Consumer Surplus(CS), Producer surplus(PS), and Dead Weight Loss(DWL)? Illustrate this in a clearly labeled graph.

b. Suppose Wenbo is able to perfectly price discriminate (also known as first degree price discrimination). What quantity will Wenbo sell? What will be his profit? What is the CS, PS, and DWL? Illustrate this in a clearly labeled graph.

c. Suppose instead Wenbo decides to practice second degree price discrimination: he charges $275 for the first 25 units, and then decreases his price by a certain amount each time he sells another 25 units. What is that amount? What is Wenbo’s profit? What is CS, PS, and DWL? Illustrate this in a clearly labeled graph.

d. Comparing (a), (b), and (c), which pricing structure do you think Wenbo prefers? Is this what you expected? Explain.
Answer #1

Soluhon : Givenノ market pe-mand p = 300-a mongiral art Re2 cost c3 ly cailb ニスas = car+ar)-(a ter) t= 112489 个 Sragh N C so PS NUR 300 So pa Mc too300 Mc9 & st しー2e ) beeoefe edg.cnnt amort 나.dlr it cs) 273 CS2 28D 200 aroeSTon: 8oD CSI CS1 Mc. 286 213 200